Line l passes through (1,-3) and is perpendicular to y = (1/5) x - 7. What is
the equation of line l ?

Answer by Tutorteddy.Com(12)

Let the equation of the line be y = mx + c, being the gradient of the line.
As the line is perpendicular to y = (1/5) x - 7,
m X (1/5) = -1
=> m = -1/1/5 = -5
Hence the required line becomes y = -5x + c
As the line passes through (1,-3)
the given equation becomes -3 = (-5) X 1 + c = -5 + c
=> c = -3 + 5 = 2
Hence the required equation of the line becomes y = -5x + 2
=> 5x + y - 2 = 0